What is the rotational kinetic energy of a 1.20-kg sphere with a radius of 0.250m rotating about its axis and an angular speed of 0.38 rad/s?
1
Expert's answer
2020-12-14T12:18:26-0500

For solid sphere (ball) of radius "r" and mass "m":


"I=(2\/5)mr^2" . So,


"KE=\frac{I\omega^2}{2}=\frac{(2/5)mr^2\omega^2}{2}="


"=\frac{mr^2\omega^2}{5}=\frac{1.2\cdot 0.25^2\cdot 0.38^2}{5}\approx0.0022(J)"